HUBBLE SPACE TELESCOPE DAILY REPORT #1865 PERIOD COVERED: 0000Z (UTC) 04/25/97 - 0000Z (UTC) 04/28/97 Daily Status Report as of 118/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of WF/PC-2 7553 (Pure Parallel Archive Survey = II) WF/PC-2 parallel observations were obtained in the ultraviolet in = search for interesting objects. This program is designed to find = extragalactic sources with near ultraviolet excesses (3000 A0) outside = the galactic plane and far ultraviolet sources (1600 A0) in the galactic = plane. The procedure was executed as scheduled, and no anomalies were = reported. 1.2 Completed Two Sets of WF/PC-2 6361 (High Resolution Imaging of = X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the = known, low, red-shifted, x-ray active, galactic nuclei (AGN) MS0516.-4 = and MS0849.5+0. These data will allow a study of the optical properties = of an unbiased sample of AGN host galaxies. The observations were = completed as scheduled, and no problems were reported. 1.3 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR06. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed NIC/1/2/3 7050 (NICMOS Differential Photometry Test) The Near Infrared Cameras (NIC One, Two and Three) were used to = make observations of standard test star P330E as part of a differential = photometric test. This is a inter-camera photometric precision and stabil= ity test, obtained from observations of the standard star at 25 = positions, in each camera. The observations were executed as scheduled, = and no anomalies were noted. 1.5 Completed NIC/1/2/3 7051 (NICMOS Detector Noise and Dark = Characterization - Non-Occulted) The Near Infrared Cameras (NIC One, Two and Three) were used to = make detector noise and dark chacturizational calibration observations. = The purpose of this proposal is to characterize NICMOS read noise and = dark current for the three cameras, and to verify the stability since = SLTV and measurements of the cosmic ray event outside SAA. The = observations were completed as planned, and no problems were reported. 1.6 Completed NIC/1/2/3 7048 (NICMOS Parallel Thermal Background Test) The Near Infrared Cameras (Cameras One, Two and Three) were used = to make observations of an object designated as Target-1 in order execute = measurements of the stability of the HST thermal emissions. The = observations were executed as scheduled, and no problems were reported. 1.7 Completed Two Sets of NIC/1/2/3 7045 (NICMOS Persistence Test) The Near Infrared Cameras (NIC One, Two and Three) were used to = make observations of standard star HD106965 in order to measure the = system's persistence. This test measures the effects of severe = over-exposure of the NICMOS detectors to determine if operational = limitations are required for programs expecting such over exposures. It = also obtained data on the extreme wings of the NICMOS PSF. Dark calibrati= on observations were also taken. At the beginning of the first iteration = of this proposal, the guide star acquisition defaulted to fine lock on = FGS 2 only when the scan search limit was exceeded. HSTAR 6045 was = written. Near the beginning of the second iteration, the acquisition = similarly defaulted, but to FGS 3. HSTAR 6047 was written. Otherwise, = the observation completed as planned, and no further problems were = reported. 1.8 Completed Two sets of WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.9 Completed STIS/CCD 7123 (MAMA and CCD Spectroscopy of SN1987A) The Space Telescope Imaging Spectrograph (CCD) was used to make = observations of the super-nova remnant LMC-SN1987 which is located in the = Large Magellanic Cloud. This is to study SN-1987A and its nebula by = taking spectra at both ultraviolet and optical wavelengths to obtain a = spatially resolved map of the debris and the circumstellar rings. A = waveline calibration image was also taken. The observations were = completed as planned, and no anomalies were reported. 1.10 Completed WF/PC-2 7029 (WF/PC-2 Cycle 6/SMOV Lyman-Alpha = Check-Part 2) The WF/PC-2 was used to make observations of WD0310-688 in order = to obtain far-ultraviolet (Lyman-Alpha) measures of the system's = throughput. This is to quantify any contamination of the WF/PC-2 = pick-off mirror caused by SMOV activities. The observations were = completed as planned, and no anomalies were reported. 1.11 Completed NIC/1/2/3 7043 (NICMOS Focus Monitor) The Near Infrared Cameras (Cameras One, Two and Three) were used = to make observations of NGC-3603 in order to make focus calibration and = monitoring measurements. This proposal is built to run NIC/1/2 focus = sweeps on a weekly basis, and NIC/3 focus sweeps twice a week during SMOV. The observations = were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6770 (Resolving the Nature of Nova Shells) The WF/PC-2 was used to make images of the post-nova remnant = BT-Mon in order to study its surrounding shells structure. Past data has = indicate a strong correlation between nova speed class and the = ellipticity of the resulting remnants. The observations were completed = as scheduled, and no problems were noted. 1.13 Completed NIC/2 7154 (NICMOS Coronographic Hole Monitor) The Near Infrared Camera (NIC 2) was used to make observations of = object NGC-3606 in order to monitor the effects of coronagraphic hole. = The purpose of this proposal is to take lamp flats with camera 2 (and = cameras 1 and 3 occasionally) to monitor the location of the = coronographic spot. The observations were completed as scheduled, and no = anomalies were reported. 1.14 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda = >From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova = SN9784 in order to study its redshift. The HST was used to make direct = measurements of the cosmological parameters Omega, Lambda, and curvature = (Omega_k) using Type 1A supernovae (SNe 1a) as a calibrated "standard = candle". The observations were executed as scheduled, and no anomalies = were noted. 1.15 Completed FGS 6878 (Parallax of NGC-6853) The Fine Guidance Sensors were used to make astrometric = observations of stars located in NGC-6853. NGC-6853 is one of the = closest planetary nebulae. Its trigonometric parallax is presently only = poorly known. HST observations have the potential to reduce the parallax = uncertainty to about 0.0005 arcsec. The observations were completed as = scheduled, and no anomalies were reported. 1.16 Completed WF/PC-2 6587 (Black Holes and Cores of Early-Type = Galaxies) The WF/PC-2 was used to make observations of early-type galaxy = IC4329 in order to study its nuclear core region and its possible black = hole. The campaign to find black holes in the centers of galaxies using = HST has so far been highly successful. The observations were completed = as planned, and no problems were reported. 1.17 Completed NIC/1/2/3 7110 (NICMOS Detector Noise and Dark = Characterization - Non Nominal) The purpose of this proposal is to characterize NICMOS read noise = and dark current for the three cameras. Verification of the stability = since SLTV and measurements of the cosmic ray event outside SAA were also = be performed. The observations were executed as scheduled, and no = anomalies were reported. 1.18 Completed WF/PC-2 6194 (WF/PC-2 Cycle 5: Polarization Proposal) The WF/PC-2 was used to make polarization observations of star = GRW+70D582 at four different spacecraft roll angles. This should permit = a detailed and independent calibration of each polarizer quad. The = observations were executed as scheduled, and no anomalies were noted. 1.19 Completed FOC/96 6426 (A Search for Multiple Images of QSOs Seen = Through Damped Ly-Alpha Absorbers) The Faint Object Camera (f/96) was used to make observations of = QSO PG0000-263. The propose of these observations is to obtain = high-resolution images of a sample of QSOs with strong damped Lyman-Alpha = absorption in their spectra. The observations should determine whether = these QSOs are multiply imaged by gravitational lensing. The = observations were completed as planned, and no problems were reported. 1.20 Completed NIC/3 7047 (NICMOS Pointed Thermal Background Test) The Near Infrared Camera (Camera Three) was used to conduct a = series of tests for the measurement of the stability of the HST thermal = emissions, through a series of pointed observations. Two hours into this = sixteen-hour proposal, at 117/0830Z, there were losses of lock that = lasted for a period of 20-minutes. HSTAR 6048 was written. The = observations was completed as planned, and affect on science was = reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 At 115/221616Z and 116/173816Z, the acquisition defaults described = in 1.7 occurred. Scheduled Reacquisitions : 20 Successful : 20 There were six to eight losses of lock as described in 1.20. 2.2 FHST Updates: Scheduled : 50 Successful : 50 2.3 Operations Notes: STIS was commanded to operate at 115/2348Z with the CCD test and = the MAMA off. The memory monitor for the STIS CCD was patched at 115/1815Z. = Other DF and NSSC-1 memory monitors were patched at 115/2047Z. There = were NSSC-1 and DF safing sequence patches uplinked at 115/2133Z. Several low voltage power supplies were turned off at 116/2041Z. STIS event flag 10 was set at 118/0115Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH